# Flock: A Knowledge Graph Foundation Model via Learning on Random Walks

### v1

@@ Appendix B Methodology - details
In this section, we expand on the descriptions of individual components of Flock summarized in Section 4: the random walk algorithm, the recording protocol, the sequence processor, and the consensus protocol.
B.1 Uniform random walk
@@
Let $G=(V,E,R)$ be a knowledge graph, and let $l$ be the length of random walks. For each node $v\in V$, we will denote by $\mathcal{N}(v)$ the set of neighbors of $v$:
@@
where $R^{-1}$ is the set symbolizing the inverses of relation types in $R$. The uniform random walk with no backtracking $\eta(G,l)$ of length $l$ over $G$, represented as:
$V_{0}\xrightarrow{R_{1}}V_{1}\xrightarrow{R_{2}}\cdots\xrightarrow{R_{l}}V_{l}$
@@
B.2 Recording function
Given a KG $G=(V,E,R)$, a query $q=(h_{q},r_{q},?)$, a walk $\bar{\eta}=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{l}}v_{l}$ of length $l$ over $G$, and a set of embeddings $\mathbf{v}$ of nodes $V$ and $\mathbf{r}$ of relations $R$, our recording function $w$ first splits the walk into a sequence of $l+1$ steps:
$(r_{0},v_{0}),(r_{1},v_{1}),\dots(r_{l},v_{l})$
@@ B.2 Recording function
$w(\bar{\eta};G,q,\mathbf{v},\mathbf{r})=(S_{0},S_{1},\dots,S_{l})$
B.3 Sequence processor
@@
$w(\bar{\eta};G,q,\mathbf{v},\mathbf{r})=(S_{0},S_{1},\dots,S_{l})$
@@
  • $\mathbf{A}_{v},\mathbf{A}_{r}\in\mathbb{R}^{(l+1)\times d}$: embedding tables for anonymized vertices and relations, respectively,
@@
$\left(\mathbf{s}^{(V)}_{i},\mathbf{s}^{(R)}_{i},\mathbf{a}^{(V)}_{i},\mathbf{a}^{(R)}_{i}\right)=\mathbf{\Omega}([\mathbf{c}_{0},\mathbf{c}_{1},\dots,\mathbf{c}_{l}])$
where $\mathbf{s}^{(V)}_{i},\mathbf{s}^{(R)}_{i}\in\mathbb{R}^{(l+1)\times h\times d_{h}}$ and $\mathbf{a}^{(V)}_{i},\mathbf{a}^{(R)}_{i}\in\mathbb{R}^{(l+1)\times h}$. Stacking all $N$ of them gives us the final output of the sequence processor.
B.4 Consensus protocol
@@
Appendix C Proofs
C.1 Expressivity
The main proposition of this section formalizes the fact that Flock can approximate any link-invariant function over fixed-size knowledge graphs in probability. Intuitively, when the length of the sampled walks $l$ becomes higher, the probability of a single walk witnessing all the edges grows to 1. Once a walk visits all the edges, a sufficiently powerful sequence processor can derive the whole graph structure from its anonymized representation, recreating the graph in its entirety, up to isomorphism. Then, the processor can return the value of the approximated function for that graph.
@@ C.1 Expressivity
Lemma C.1.
Let $G\in\mathbb{K}_{n,m}$ for some $n,m$. The edge cover time $C_{E}(G)$ of $G$, using the algorithm from Section B.1, is finite.
@@
Let $G=(V,E,R)\in\mathbb{K}_{n,m}$ be a graph. For any edge $e\in E$ and any vertex $v\in V$, let $H_{v}(e)$ denote the expected number of steps of the random walk algorithm $\eta$ described in Section B.1. Then, the edge cover time $C_{E}(G)$ of $G$ with $\eta$, i.e. the expected number of steps that $\eta$ needs to take before visiting every edge in $G$, is bounded above by:
@@
Moreover, let $v_{i(d+2)}=u_{0}\xrightarrow{s_{1}}u_{1}\xrightarrow{s_{2}}\dots\xrightarrow{s_{l}}u_{l}\in\{h_{e},t_{e}\}$ be a shortest path from $v_{i(d+2)}$ to one of $h_{e},t_{e}$. Note that by minimality, there cannot be any backtracking while following this path. Therefore, the probability of the next visited node is dependent only on the value of the previous one, and we can bound the probability $P(A_{i})$ of $A_{i}$ from below by:
${\mathbb{P}}(A_{i})\geq{\mathbb{P}}(\text{pass through }e\text{ after reaching }h_{e}\text{ or }t_{e})\cdot\prod_{j=0}^{l-1}{\mathbb{P}}(v_{i(d+2)+j+1}=u_{j+1}\mid v_{i(d+2)+j}=u_{j})$
@@
${\mathbb{P}}(A_{i})$ $\geq{\mathbb{P}}(\text{pass through }e\text{ after reaching }h_{e}\text{ or }t_{e})\cdot\prod_{j=0}^{l-1}{\mathbb{P}}(v_{i(d+2)+j+1}=u_{j+1}\mid v_{i(d+2)+j}=u_{j})$
$\geq\frac{1}{\Delta\cdot\rho}\cdot\prod_{j=0}^{l-1}\frac{1}{\Delta}\quad\geq\quad\frac{1}{\Delta\cdot\rho}\left(\frac{1}{\Delta}\right)^{l}\quad\geq\quad\frac{1}{\Delta\cdot\rho}\left(\frac{1}{\Delta}\right)^{d}\quad=\quad\frac{1}{\rho\Delta^{d+1}}$
@@
Remark C.2.
The bound obtained in the proof of Lemma C.1 is very crude. In fact, we could transform the given knowledge graph into a simple graph (undirected, with no multi-edges) by substituting each edge $u\xrightarrow{r}v$ with two undirected edges $u\leftrightarrow v_{(u,r,v)}\leftrightarrow v$. The augmented graph will then have $n+m$ vertices, and our random walk algorithm naturally translates to a weighted random walk on the transformed graph. This hints at an assumption that in practice, the edge cover time of the used random walk algorithm is of the magnitude $O((n+m)^{3})=O(n^{3}+m^{3})$.
@@
Lemma C.4.
For each pair $(n,m)$, there exists a number $C_{n,m}$ such that the edge cover time, using the algorithm from Section B.1, of any knowledge graph in $\mathbb{K}_{n,m}$ is at most $C_{n,m}$.
@@
The result follows from Lemmas C.1 and C.3. As two isomorphic graphs have identical cover time, we can set $C_{n,m}$ to be the maximum of cover times of representatives of all isomorphic classes, which, by finiteness of both, is well-defined. ∎
Lemma C.5.
Let $G\in\mathbb{K}_{n,m}$ be a graph, $q=(h_{q},r_{q},?)$ be a link query over $G$, and $\bar{\eta}$ be a walk over $G$. If $\bar{\eta}$ traverses all edges of $G$, then using only the output $w(\bar{\eta};G,q,\cdot,\cdot)$ of the recording function $w$ detailed in Section B.2, we can construct a graph-query pair $(H,q^{\prime})$ isomorphic to $(G,q)$.
@@
Suppose that $\bar{\eta}=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{l}}v_{l}$ visits all edges of $G=(V,E,R)$ and let $l$ be its length. Recall the anonymization functions $\operatorname{id}_{V}(\cdot;\bar{\eta})$ and $\operatorname{id}_{R}(\cdot;\bar{\eta})$ as defined in Section B.2. The output $w(\bar{\eta};G,q,\cdot,\cdot)$ (the embedding functions provided as the last two arguments are irrelevant) is a sequence of tuples $S_{0},S_{1},\dots,S_{l}$ with each $S_{i}$ equal to:
@@
Then by the definition of $w$ (Section B.2), it is straightforward to check that the pair $(\operatorname{id}_{V}(\cdot;\bar{\eta}),\operatorname{id}_{R}(\cdot;\bar{\eta}))$ defines an isomorphism from $(G,q)$ to $(H,q^{\prime})$. Indeed, both these functions are injective by construction, and as $\bar{\eta}$ witnesses all nodes and relations, they are well-defined bijections. For each unique edge traversed by $\bar{\eta}$, there exists a unique edge in $E^{\prime}$ translated to the anonymized space, which implies an isomorphism between $E$ and $E^{\prime}$. Finally, by utilizing the flags $\delta_{v_{i}=h_{q}}$ and $\delta_{r_{j}=r_{q}}$, we can identify the query head node and relation in the new graph. All things considered, we can reconstruct the pair $(G,q)$, up to isomorphism, from the output of $w(\bar{\eta};G,q,\cdot,\cdot)$. ∎
We are now ready to prove the main result regarding the universality of Flock as an approximation of link invariant functions. The outline of the proof is as follows: 1) Using the upper-bound on the edge cover time of graphs in $\mathbb{K}_{n,m}$ derived in Lemma C.4, we can bound the probability of sampling a walk that visits all edges, 2) Once such a walk is sampled, we can recover the graph and query, up to isomorphism, from its anonymized form (Lemma C.5), 3) Lastly, we can return the value of the approximated function for the derived isomorphic instance. Since the approximated function is link invariant, if the reconstructed graph matches the original one, we return precisely the correct value.
@@
With a powerful enough sequence processor $f_{\theta}$, the Flock framework described in Section 4 is a universal approximator of link invariant functions over $\mathbb{K}_{n,m}$ for all pairs $(n,m)$.
@@
For simplicity, let us consider a situation where only a single walk $\bar{\eta}$ of length $l$ is sampled by the model (otherwise, omit additional walks). We will also restrict the argument to a single refinement case – the result can be extended to multiple refinement steps by returning $\Delta\mathbf{v},\Delta\mathbf{r}=0$ during all additional iterations. Consider a sequence processor $f_{\theta}$ that given the output $w(\bar{\eta};G,q,\cdot,\cdot)$ of the recording protocol, creates a graph-query pair $(H,q^{\prime})$ with $q^{\prime}=(h_{q^{\prime}},r_{q^{\prime}},?)$ using the strategy described in the proof of Lemma C.5, and returns a vector $\mathbf{h}\in\mathbb{R}^{l+1}$ whose $i^{\text{th}}$ entry is equal $\mathbf{h}_{i}=\varphi(H)((h_{q^{\prime}},r_{q^{\prime}},\operatorname{id}_{V}(v_{i};\bar{\eta}))$ where $v_{i}$ is the $i^{\text{th}}$ node visited by $\bar{\eta}$. The consensus protocol $c$, provided $t$ was visited by $\bar{\eta}$, can then identify $t$ as one of the $v_{j}$ and pull the corresponding embedding $\mathbf{h}_{j}=\varphi(H)((h_{q^{\prime}},r_{q^{\prime}},\operatorname{id}_{V}(t;\bar{\eta}))$, returning it as the output $\mathbf{v}(t)=\mathbf{h}_{j}$ (note that no matter which specific value of $j$ is chosen, this value will be the same). Finally, the classification head can work as an identity operation, returning $X_{\theta}(G,q)(t)=\mathbf{v}(t)=\varphi(H)((h_{q^{\prime}},r_{q^{\prime}},\operatorname{id}_{V}(t;\bar{\eta}))$.
@@
By Lemma C.5, in such case, the reconstructed pair $(H,q^{\prime})$ is isomorphic to $(G,q)$ by the isomorphism $id=(\operatorname{id}_{V}(\cdot;\bar{\eta}),\operatorname{id}_{R}(\cdot;\bar{\eta}))$. Since $\varphi$ is link invariant, we can write:
@@
Hence, to show (9), it suffices to prove that we can uniformly choose the length $l$ of the random walk so that the probability of $\bar{\eta}$ covering all the edges is greater than $1-\delta$. By Markov’s inequality:
@@
$=\frac{C_{E}(G)}{l}$
But by Lemma C.4, $C_{E}(G)\leq C_{n,m}$ for some constant $C_{n,m}$. Hence, taking $l>\frac{C_{n,m}}{\delta}$, we get:
${\mathbb{P}}(\bar{\eta}\text{ does not cover all edges})\leq\frac{C_{E}(G)}{l}\leq\frac{C_{n,m}}{l}<\delta$
This means that for such a choice of $l$:
@@
which leads to the conclusion that for $l>\frac{C_{n,m}}{\delta}$, the proposed Flock framework satisfies:

### v2

@@ Appendix A Methodology - details
In this section, we expand on the descriptions of individual components of Flock summarized in Section˜4: the random walk algorithm, the recording protocol, the sequence processor, and the consensus protocol.
A.1 Uniform random walk
@@
Let $G=(V,E,R)$ be a knowledge graph, and let $\ell$ be the length of random walks. For each node $v\in V$, we will denote by $\mathcal{N}(v)$ the set of neighbors of $v$:
@@
where $R^{-1}$ is the set symbolizing the inverses of relation types in $R$. The uniform random walk with no backtracking $\eta(G,\ell)$ of length $\ell$ over $G$, represented as:
$V_{0}\xrightarrow{R_{1}}V_{1}\xrightarrow{R_{2}}\cdots\xrightarrow{R_{\ell}}V_{\ell}$
@@
A.2 Recording function
Given a KG $G=(V,E,R)$, a query $q=(h_{q},r_{q},?)$, a walk $\bar{\eta}=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{\ell}}v_{\ell}$ of length $\ell$ over $G$, and a set of embeddings $\mathbf{v}$ of nodes $V$ and $\mathbf{r}$ of relations $R$, our recording function $w$ first splits the walk into a sequence of $\ell+1$ steps:
$(r_{0},v_{0}),(r_{1},v_{1}),\dots(r_{\ell},v_{\ell})$
@@ A.2 Recording function
$w(\bar{\eta};G,q,\mathbf{v},\mathbf{r})=(S_{0},S_{1},\dots,S_{\ell})$
A.3 Sequence processor
@@
$w(\bar{\eta};G,q,\mathbf{v},\mathbf{r})=(S_{0},S_{1},\dots,S_{\ell})$
@@
  • $\mathbf{A}_{v},\mathbf{A}_{r}\in\mathbb{R}^{(\ell+1)\times d}$: embedding tables for anonymized vertices and relations, respectively,
@@
$\left(\mathbf{s}^{(V)}_{i},\mathbf{s}^{(R)}_{i},\mathbf{a}^{(V)}_{i},\mathbf{a}^{(R)}_{i}\right)=\mathbf{\Omega}([\mathbf{c}_{0},\mathbf{c}_{1},\dots,\mathbf{c}_{\ell}])$
where $\mathbf{s}^{(V)}_{i},\mathbf{s}^{(R)}_{i}\in\mathbb{R}^{(\ell+1)\times h\times d_{h}}$ and $\mathbf{a}^{(V)}_{i},\mathbf{a}^{(R)}_{i}\in\mathbb{R}^{(\ell+1)\times h}$. Stacking all $N$ of them gives us the final output of the sequence processor.
A.4 Consensus protocol
@@
Appendix B Proofs
B.1 Expressivity
The main proposition of this section formalizes the fact that Flock can approximate any link-invariant function over fixed-size knowledge graphs in probability. Intuitively, when the length of the sampled walks $\ell$ becomes higher, the probability of a single walk witnessing all the edges grows to 1. Once a walk visits all the edges, a sufficiently powerful sequence processor can derive the whole graph structure from its anonymized representation, recreating the graph in its entirety, up to isomorphism. Then, the processor can return the value of the approximated function for that graph.
@@ B.1 Expressivity
Lemma B.1.
Let $G\in\mathbb{K}_{n,m}$ for some $n,m$. The edge cover time $C_{E}(G)$ of $G$, using the algorithm from Section˜A.1, is finite.
@@
Let $G=(V,E,R)\in\mathbb{K}_{n,m}$ be a graph. For any edge $e\in E$ and any vertex $v\in V$, let $H_{v}(e)$ denote the expected number of steps of the random walk algorithm $\eta$ described in Section˜A.1. Then, the edge cover time $C_{E}(G)$ of $G$ with $\eta$, i.e. the expected number of steps that $\eta$ needs to take before visiting every edge in $G$, is bounded above by:
@@
Moreover, let $v_{i(d+2)}=u_{0}\xrightarrow{s_{1}}u_{1}\xrightarrow{s_{2}}\dots\xrightarrow{s_{\ell}}u_{\ell}\in\{h_{e},t_{e}\}$ be a shortest path from $v_{i(d+2)}$ to one of $h_{e},t_{e}$. Note that by minimality, there cannot be any backtracking while following this path. Therefore, the probability of the next visited node is dependent only on the value of the previous one, and we can bound the probability $P(A_{i})$ of $A_{i}$ from below by:
${\mathbb{P}}(A_{i})\geq{\mathbb{P}}(\text{pass through }e\text{ after reaching }h_{e}\text{ or }t_{e})\cdot\prod_{j=0}^{\ell-1}{\mathbb{P}}(v_{i(d+2)+j+1}=u_{j+1}\mid v_{i(d+2)+j}=u_{j})$
@@
${\mathbb{P}}(A_{i})$ $\geq{\mathbb{P}}(\text{pass through }e\text{ after reaching }h_{e}\text{ or }t_{e})\cdot\prod_{j=0}^{\ell-1}{\mathbb{P}}(v_{i(d+2)+j+1}=u_{j+1}\mid v_{i(d+2)+j}=u_{j})$
$\geq\frac{1}{\Delta\cdot\rho}\cdot\prod_{j=0}^{\ell-1}\frac{1}{\Delta}\quad\geq\quad\frac{1}{\Delta\cdot\rho}\left(\frac{1}{\Delta}\right)^{\ell}\quad\geq\quad\frac{1}{\Delta\cdot\rho}\left(\frac{1}{\Delta}\right)^{d}\quad=\quad\frac{1}{\rho\Delta^{d+1}}$
@@
Remark B.2.
The bound obtained in the proof of Lemma B.1 is very crude. In fact, we could transform the given knowledge graph into a simple graph (undirected, with no multi-edges) by substituting each edge $u\xrightarrow{r}v$ with two undirected edges $u\leftrightarrow v_{(u,r,v)}\leftrightarrow v$. The augmented graph will then have $n+m$ vertices, and our random walk algorithm naturally translates to a weighted random walk on the transformed graph. This hints at an assumption that in practice, the edge cover time of the used random walk algorithm is of the magnitude $O((n+m)^{3})=O(n^{3}+m^{3})$.
@@
Lemma B.4.
For each pair $(n,m)$, there exists a number $C_{n,m}$ such that the edge cover time, using the algorithm from Section˜A.1, of any knowledge graph in $\mathbb{K}_{n,m}$ is at most $C_{n,m}$.
@@
The result follows from Lemmas B.1 and B.3. As two isomorphic graphs have identical cover time, we can set $C_{n,m}$ to be the maximum of cover times of representatives of all isomorphic classes, which, by finiteness of both, is well-defined. ∎
Lemma B.5.
Let $G\in\mathbb{K}_{n,m}$ be a graph, $q=(h_{q},r_{q},?)$ be a link query over $G$, and $\bar{\eta}$ be a walk over $G$. If $\bar{\eta}$ traverses all edges of $G$, then using only the output $w(\bar{\eta};G,q,\cdot,\cdot)$ of the recording function $w$ detailed in Section˜A.2, we can construct a graph-query pair $(H,q^{\prime})$ isomorphic to $(G,q)$.
@@
Suppose that $\bar{\eta}=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{\ell}}v_{\ell}$ visits all edges of $G=(V,E,R)$ and let $\ell$ be its length. Recall the anonymization functions $\operatorname{id}_{V}(\cdot;\bar{\eta})$ and $\operatorname{id}_{R}(\cdot;\bar{\eta})$ as defined in Section˜A.2. The output $w(\bar{\eta};G,q,\cdot,\cdot)$ (the embedding functions provided as the last two arguments are irrelevant) is a sequence of tuples $S_{0},S_{1},\dots,S_{\ell}$ with each $S_{i}$ equal to:
@@
Then by the definition of $w$ (Section˜A.2), it is straightforward to check that the pair $(\operatorname{id}_{V}(\cdot;\bar{\eta}),\operatorname{id}_{R}(\cdot;\bar{\eta}))$ defines an isomorphism from $(G,q)$ to $(H,q^{\prime})$. Indeed, both these functions are injective by construction, and as $\bar{\eta}$ witnesses all nodes and relations, they are well-defined bijections. For each unique edge traversed by $\bar{\eta}$, there exists a unique edge in $E^{\prime}$ translated to the anonymized space, which implies an isomorphism between $E$ and $E^{\prime}$. Finally, by utilizing the flags $\delta_{v_{i}=h_{q}}$ and $\delta_{r_{j}=r_{q}}$, we can identify the query head node and relation in the new graph. All things considered, we can reconstruct the pair $(G,q)$, up to isomorphism, from the output of $w(\bar{\eta};G,q,\cdot,\cdot)$. ∎
We are now ready to prove the main result regarding the universality of Flock as an approximation of link invariant functions. The outline of the proof is as follows: 1) Using the upper-bound on the edge cover time of graphs in $\mathbb{K}_{n,m}$ derived in Lemma B.4, we can bound the probability of sampling a walk that visits all edges, 2) Once such a walk is sampled, we can recover the graph and query, up to isomorphism, from its anonymized form (Lemma B.5), 3) Lastly, we can return the value of the approximated function for the derived isomorphic instance. Since the approximated function is link invariant, if the reconstructed graph matches the original one, we return precisely the correct value.
@@
With a powerful enough sequence processor $f_{\theta}$, the Flock framework described in Section˜4 is a universal approximator of link invariant functions over $\mathbb{K}_{n,m}$ for all pairs $(n,m)$.
@@
For simplicity, let us consider a situation where only a single walk $\bar{\eta}$ of length $\ell$ is sampled by the model (otherwise, omit additional walks). We will also restrict the argument to a single refinement case – the result can be extended to multiple refinement steps by returning $\Delta\mathbf{v},\Delta\mathbf{r}=0$ during all additional iterations. Consider a sequence processor $f_{\theta}$ that given the output $w(\bar{\eta};G,q,\cdot,\cdot)$ of the recording protocol, creates a graph-query pair $(H,q^{\prime})$ with $q^{\prime}=(h_{q^{\prime}},r_{q^{\prime}},?)$ using the strategy described in the proof of Lemma B.5, and returns a vector $\mathbf{h}\in\mathbb{R}^{l+1}$ whose $i^{\text{th}}$ entry is equal $\mathbf{h}_{i}=\varphi(H)((h_{q^{\prime}},r_{q^{\prime}},\operatorname{id}_{V}(v_{i};\bar{\eta}))$ where $v_{i}$ is the $i^{\text{th}}$ node visited by $\bar{\eta}$. The consensus protocol $c$, provided $t$ was visited by $\bar{\eta}$, can then identify $t$ as one of the $v_{j}$ and pull the corresponding embedding $\mathbf{h}_{j}=\varphi(H)((h_{q^{\prime}},r_{q^{\prime}},\operatorname{id}_{V}(t;\bar{\eta}))$, returning it as the output $\mathbf{v}(t)=\mathbf{h}_{j}$ (note that no matter which specific value of $j$ is chosen, this value will be the same). Finally, the classification head can work as an identity operation, returning $X_{\theta}(G,q)(t)=\mathbf{v}(t)=\varphi(H)((h_{q^{\prime}},r_{q^{\prime}},\operatorname{id}_{V}(t;\bar{\eta}))$.
@@
By Lemma B.5, in such case, the reconstructed pair $(H,q^{\prime})$ is isomorphic to $(G,q)$ by the isomorphism $id=(\operatorname{id}_{V}(\cdot;\bar{\eta}),\operatorname{id}_{R}(\cdot;\bar{\eta}))$. Since $\varphi$ is link invariant, we can write:
@@
Hence, to show (4), it suffices to prove that we can uniformly choose the length $\ell$ of the random walk so that the probability of $\bar{\eta}$ covering all the edges is greater than $1-\delta$. By Markov’s inequality:
@@
$=\frac{C_{E}(G)}{\ell}$
But by Lemma B.4, $C_{E}(G)\leq C_{n,m}$ for some constant $C_{n,m}$. Hence, taking $\ell>\frac{C_{n,m}}{\delta}$, we get:
${\mathbb{P}}(\bar{\eta}\text{ does not cover all edges})\leq\frac{C_{E}(G)}{\ell}\leq\frac{C_{n,m}}{\ell}<\delta$
This means that for such a choice of $\ell$:
@@
which leads to the conclusion that for $\ell>\frac{C_{n,m}}{\delta}$, the proposed Flock framework satisfies: